# FINITE-DIMENSIONAL GAUSSIAN APPROXIMATION FOR DEEP NEURAL NETWORKS: UNIVERSALITY IN RANDOM WEIGHTS

### v1

@@ 2.1 Error for general to Gaussian weights
For the first term of (2.2), we are comparing $W^{(L-1)}\sigma(F^{(L-1)})$ and $\widetilde{W}^{(L-1)}\sigma(F^{(L-1)})$, and, in anticipation of conditioning on $F^{(L-1)}$, we state and prove the following lemma bounding the $\mathsf{d}_{3}$ distance in this setting when thinking of $\sigma(F^{(L-1)})$ as fixed.
@@
Let $h\in\mathbbm{R}^{m\times s}$ and $W$ be a $d\times m$ random matrix which has centered independent entries having the same variance $\mathop{\mathrm{Var}}\nolimits(W_{ij})=:c_{2}/m$, also satisfying $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert W_{ij}\rvert^{3}\bigr{]}\leqslant c_{3}/m^{3/2}$. Define $F\in\mathbbm{R}^{d\times s}$ by
@@
Analogously, let $\widetilde{W}$ be a $d\times m$ random matrix that has i.i.d. centered Gaussian entries with the same variances as $W_{ij}$, and define $\widetilde{F}\in\mathbbm{R}^{n\times s}$ by
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\widetilde{F}_{i,a},\widetilde{F}_{j,b}\bigr{]}=\sum_{k,l=1}^{m}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\widetilde{W}_{ik}\widetilde{W}_{jl}\bigr{]}h_{k,a}h_{l,b}=\delta_{ij}c_{2}\Biggl{(}\frac{1}{m}\sum_{k=1}^{m}h_{k,a}h_{k,b}\Biggr{)}=:\delta_{ij}\widetilde{C}_{a,b}.$
@@
$\bigl{[}\mathcal{A}_{\widetilde{C}}\,\eta(F)\bigr{]}:=\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a,b=1}^{s}\widetilde{C}_{a,b}\,\partial_{(i,a),(i,b)}\eta(F)\biggr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a=1}^{s}F_{i,a}\partial_{(i,a)}\eta(F)\biggr{]},$ (2.3)
@@
Working on
$\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a=1}^{s}F_{i,a}\partial_{(i,a)}\eta(F)\biggr{]}=\sum_{i=1}^{d}\sum_{a=1}^{s}\sum_{k=1}^{m}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}W_{ik}h_{k,a}\partial_{(i,a)}\eta(F)\bigr{]},$ (2.5)
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}W_{ik}h_{k,a}\partial_{(i,a)}\eta(F)\bigr{]}=\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}W_{ik}h_{k,a}\bigl{(}\partial_{(i,a)}\eta(F)-\partial_{(i,a)}\eta(F-V_{ik})\bigr{)}\bigr{]}.$ (2.6)
@@
Plugging this in (2.6) and then (2.5), and using that $\mathop{{}\mathbbm{E}}\mathopen{}[W_{ik}^{2}]=c_{w}/m$ and the independence of $W_{ik}$ and $F-V_{ik}$, we can bound the absolute value of (2.3) as
[…]
$\leqslant\biggl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a,b=1}^{s}\widetilde{C}_{a,b}\bigl{(}\partial_{(i,a),(i,b)}\eta(F)-\partial_{(i,a),(i,b)}\eta(F-V_{ik})\bigr{)}\biggr{]}\biggr{\rvert}$
$+\biggl{\lvert}\sum_{i=1}^{d}\sum_{a,b=1}^{s}\sum_{k=1}^{m}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}W_{ik}^{2}h_{k,a}h_{k,b}\int_{0}^{1}\bigl{(}\partial_{(i,a),(i,b)}\eta(F-V_{ik}+sV_{ik})-\partial_{(i,a),(i,b)}\eta(F-V_{ik})\bigr{)}\mathrm{Leb}(ds)\bigr{]}\biggr{\rvert}$
$\leqslant\frac{1}{3}\sum_{i=1}^{d}\sum_{a,b,c=1}^{s}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{\lvert}\widetilde{C}_{a,b}h_{k,c}W_{ik}\bigr{\rvert}+\frac{1}{6}\sum_{i=1}^{d}\sum_{a,b,c=1}^{s}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{\lvert}W_{ik}^{3}h_{k,a}h_{k,b}h_{k,c}\bigr{\rvert}$
@@
where in the second to last inequality we use the bounds on the derivatives given at (2.4) (for both terms), and the final inequality uses Cauchy-Schwarz with $\mathop{{}\mathbbm{E}}\mathopen{}[W_{ik}^{2}]=c_{2}/m$ and that $\mathop{{}\mathbbm{E}}\mathopen{}[\lvert W_{ik}\rvert^{3}]=c_{3}/m^{3/2}$. The result now easily follows. ∎
Now, conditioning on $F^{(L-1)}$, using that $F^{(L-1)}_{i}\stackrel{{\scriptstyle d}}{{=}}F^{(L-1)}_{j}$ and the monotonicity of $p$-norms in $\mathbbm{R}^{s}$, Lemma 2.1 with $h_{k,a}:=\sigma(F_{k}^{(L-1)}(x_{a}))$ gives the following corollary bounding the first term of (2.2).
@@
Retaining the notation above, if $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert W_{ij}^{(L-1)}\rvert^{3}\bigr{]}\leqslant c_{3}^{(L-1)}n_{L-1}^{-3/2}$, then
$\mathsf{d}_{3}(F_{[d]}^{(L)},\widetilde{F}_{[d]}^{(L)})\leqslant\biggl{(}\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr{)}\frac{d}{\sqrt{n_{L-1}}}\sum_{a=1}^{s}\mathop{{}\mathbbm{E}}\mathopen{}\Bigl{[}\bigl{\lvert}\sigma(F_{1}^{(L-1)}(x_{a})\bigr{\rvert}^{3}\Bigr{]}.$ (2.7)
@@ 2.2 Error for Gaussian weights to Gaussian
Let $H\in\mathbbm{R}^{m\times s}$ be a random matrix with exchangeable rows, independent of the matrix $\widetilde{W}\in\mathbbm{R}^{d\times m}$, which has i.i.d. centered independent Gaussian entries having variance $\mathop{\mathrm{Var}}\nolimits(\widetilde{W}_{ij})=:c_{2}/m$. Define $\widetilde{F}\in\mathbbm{R}^{d\times s}$ by
@@
$\mathop{{}\mathbbm{E}}\mathopen{}[G_{i,a}G_{j,b}]=\delta_{ij}c_{2}\mathop{{}\mathbbm{E}}\mathopen{}[K_{a}K_{b}],$
@@
$\mathsf{d}_{3}(\widetilde{F},G)\leqslant\frac{c_{2}\,d}{2}\sum_{a,b=1}^{s}\Bigl{(}\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}[K_{a}K_{b}]-\mathop{{}\mathbbm{E}}\mathopen{}[H_{1,a}H_{1,b}]\bigr{\rvert}+\sqrt{\frac{\mathop{\mathrm{Var}}\nolimits(H_{1,a}H_{1,b})}{m}}+\sqrt{\mathrm{Cov}(H_{1,a}H_{1,b},H_{2,a}H_{2,b})}\Bigr{)}.$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\mathcal{A}\,\eta(\widetilde{F})\bigr{]}:=\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a,b=1}^{s}c_{2}\mathop{{}\mathbbm{E}}\mathopen{}[K_{a}K_{b}]\,\partial_{(i,a),(i,b)}\eta(\widetilde{F})\biggr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a=1}^{s}\widetilde{F}_{i,a}\partial_{(i,a)}\eta(\widetilde{F})\biggr{]},$ (2.8)
@@
$\mathop{{}\mathbbm{E}}\mathopen{}[F_{i,a}F_{j,b}|H]=\delta_{i,j}\frac{c_{2}}{m}\sum_{k=1}^{m}H_{k,a}H_{k,b}=:\delta_{i,j}C_{a,b}(H).$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a=1}^{s}\widetilde{F}_{i,a}\partial_{(i,a)}\eta(\widetilde{F})\Big{|}H\biggr{]}=\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\sum_{i=1}^{d}\sum_{a,b=1}^{s}C_{a,b}(H)\,\partial_{(i,a),(i,b)}\eta(\widetilde{F})\Big{|}H\biggr{]}.$ (2.10)
@@
$\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\mathcal{A}\,\eta(\widetilde{F})\bigr{]}\bigr{\rvert}\leqslant\frac{1}{2}\sum_{i=1}^{d}\sum_{a,b=1}^{s}\Bigl{(}\bigl{\lvert}c_{2}\mathop{{}\mathbbm{E}}\mathopen{}[K_{a}K_{b}]-\mathop{{}\mathbbm{E}}\mathopen{}[C_{a,b}(H)]\bigr{\rvert}+\mathop{{}\mathbbm{E}}\mathopen{}\bigl{\lvert}C_{a,b}(H)-\mathop{{}\mathbbm{E}}\mathopen{}[C_{a,b}(H)]\bigr{\rvert}\Bigr{)}.$
Using the exchangeability of the rows of $H$, we easily find
$\mathop{{}\mathbbm{E}}\mathopen{}[C_{a,b}(H)]=c_{2}\mathop{{}\mathbbm{E}}\mathopen{}[H_{1,a}H_{1,b}],$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\lvert C_{a,b}(H)-\mathop{{}\mathbbm{E}}\mathopen{}[C_{a,b}(H)]\rvert]$ $\leqslant\sqrt{\mathop{\mathrm{Var}}\nolimits(C_{a,b}(H))}$
@@
Taking $H_{k,a}=\sigma(F_{k}^{(L-1)}(x_{a}))$ and $K_{a}=\sigma(G_{1}^{(L-1)}(x_{a}))$ in Lemma 2.4 easily gives the following corollary bounding the second term of (2.2).
@@
$\mathsf{d}_{3}($ $\widetilde{F}_{[d]}^{(L)},G_{[d]}^{(L)})$
$\leqslant\frac{c_{2}^{(L-1)}\,d}{2}\sum_{a,b=1}^{s}\biggl{\{}\Bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\bigr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr{]}\Bigr{\rvert}$ (2.11)
@@
Let $X,Z\in\mathbbm{R}^{d}$ be random vectors with $\mathop{{}\mathbbm{E}}\mathopen{}[\lvert X_{i}\rvert^{q+d}]<C_{q}$ and $\mathop{{}\mathbbm{E}}\mathopen{}[\lvert Z_{i}\rvert^{q+d}]<C_{q}$ for $i=1,\ldots,d$ and some $q>0$. Then
$\biggl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\prod_{i=1}^{d}X_{i}-\mathop{{}\mathbbm{E}}\mathopen{}\prod_{i=1}^{d}Z_{i}\biggr{\rvert}\leqslant 2\sqrt{d}(4\sqrt{d}\,C_{q})^{(d-1)/(q+d-1)}\mathsf{d}_{1}(X,Z)^{q/(q+d-1)}.$
@@
For any coupling $(X,Z)$ of $X$ and $Z$, writing the product as a telescoping sum gives
$\biggl{\lvert}\prod_{i=1}^{d}X_{i}-\prod_{i=1}^{d}Z_{i}\biggr{\rvert}$ $\leqslant\sum_{j=1}^{d}\lvert X_{j}-Z_{j}\rvert\prod_{i=1}^{j-1}\lvert Z_{i}\rvert\prod_{k=j+1}^{d}\lvert X_{k}\rvert$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\biggl{\{}\prod_{i=1}^{d}\lvert X_{i}\rvert\mathbbm{I}\bigl{[}\lvert Z_{j}\rvert\geqslant\mathrm{T}\bigr{]}\biggr{\}}$ $\leqslant\mathrm{T}^{-q}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{\{}\lvert Z_{j}\rvert^{q}\prod_{i=1}^{d}\lvert X_{i}\rvert\biggr{\}}$
$\leqslant\mathrm{T}^{-q}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert Z_{j}\rvert^{q+d}\bigr{]}^{q/(q+d)}\prod_{i=1}^{d}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert X_{i}\rvert^{q+d}\bigr{]}^{1/q+d}$
@@
$\biggl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\prod_{i=1}^{d}X_{i}-\mathop{{}\mathbbm{E}}\mathopen{}\prod_{i=1}^{d}Z_{i}\biggr{\rvert}\leqslant\sqrt{d}\mathrm{T}^{d-1}\mathsf{d}_{1}(X,Z)+\mathrm{T}^{-q}(4d\,C_{q}).$
@@
Fix $p\in\mathbbm{N}$. If for $\ell=0,\ldots,L-2$, there are constants $c_{2p}^{(\ell)}\geqslant 1$ such that $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(W_{ij}^{(\ell)})^{2p}\bigr{]}\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, then
$\max\Bigl{\{}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]},\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(G_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]}\Bigr{\}}\leqslant B_{2p}^{(L-1)}(x),$
@@
Comparing to $\sigma(0)$ and using the triangle inequality for the $(2p)$-moment norm, we find
$\begin{split}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]}^{1/(2p)}&=\mathop{{}\mathbbm{E}}\mathopen{}\Bigl{[}\bigl{(}\sigma(F_{1}^{(L-1)}(x))-\sigma(0)+\sigma(0)\bigr{)}^{2p}\Bigr{]}^{1/(2p)}\\ &\leqslant\mathrm{Lip}_{\sigma}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(F_{1}^{(L-1)}(x))^{2p}\bigr{]}^{1/(2p)}+\lvert\sigma(0)\rvert.\end{split}$ (2.15)

### v2

@@ 2.1 Error for general to Gaussian weights
For the first term of (2.2), we are comparing $W^{(L-1)}\sigma(F^{(L-1)}(\chi))$ and $\widetilde{W}^{(L-1)}\sigma(F^{(L-1)}(\chi))$, and, in anticipation of conditioning on $F^{(L-1)}$, we state and prove the following lemma bounding the $\mathsf{d}_{3}$ distance in this setting when thinking of $\sigma(F^{(L-1)})$ as fixed.
@@
Let $h\in\mathbbm{R}^{m\times s}$ and $W$ be a $d\times m$ random matrix which has centered independent entries having the same variance $\mathop{\mathrm{Var}}\nolimits(W_{ij})=:c_{2}/m$, also satisfying $\mathbbm{E}\bigl[\lvert W_{ij}\rvert^{3}\bigr]\leqslant c_{3}/m^{3/2}$. Define $F\in\mathbbm{R}^{d\times s}$ by
@@
Analogously, let $\widetilde{W}$ be a $d\times m$ random matrix that has independent centered Gaussian entries with the same variances as $W_{ij}$, and define $\widetilde{F}\in\mathbbm{R}^{d\times s}$ by
@@
$\mathbbm{E}\bigl[\widetilde{F}_{i,a},\widetilde{F}_{j,b}\bigr]=\sum_{k,l=1}^{m}\mathbbm{E}\bigl[\widetilde{W}_{ik}\widetilde{W}_{jl}\bigr]h_{k,a}h_{l,b}=\delta_{ij}c_{2}\Biggl(\frac{1}{m}\sum_{k=1}^{m}h_{k,a}h_{k,b}\Biggr)=:\delta_{ij}\widetilde{C}_{a,b}.$
@@
$\mathbbm{E}\bigl[\mathcal{A}_{\widetilde{C}}\,\eta(F)\bigr]:=\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a,b=1}^{s}\widetilde{C}_{a,b}\,\partial_{(i,a),(i,b)}\eta(F)\biggr]-\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a=1}^{s}F_{i,a}\partial_{(i,a)}\eta(F)\biggr],$ (2.3)
@@
$\lvert\partial_{\alpha}\eta(f)\rvert\leqslant 1,\ \ \lvert\partial_{\alpha,\beta}\eta(f)\rvert\leqslant 1/2,\,\text{ and }\,\lvert\partial_{\alpha,\beta,\gamma}\eta(f)\rvert\leqslant 1/3,$ (2.4)
for all $f$. Working on
$\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a=1}^{s}F_{i,a}\partial_{(i,a)}\eta(F)\biggr]=\sum_{i=1}^{d}\sum_{a=1}^{s}\sum_{k=1}^{m}\mathbbm{E}\bigl[W_{ik}h_{k,a}\partial_{(i,a)}\eta(F)\bigr],$ (2.5)
@@
$\mathbbm{E}\bigl[W_{ik}h_{k,a}\partial_{(i,a)}\eta(F)\bigr]=\mathbbm{E}\bigl[W_{ik}h_{k,a}\bigl(\partial_{(i,a)}\eta(F)-\partial_{(i,a)}\eta(F-V_{ik})\bigr)\bigr].$ (2.6)
@@
Plugging this in (2.6) and then (2.5), and using that $\mathbbm{E}[W_{ik}^{2}]=c_{2}/m$ and the independence of $W_{ik}$ and $F-V_{ik}$, we can bound the absolute value of (2.3) as
$\bigl\lvert$ $\mathbbm{E}\bigl[\mathcal{A}_{\widetilde{C}}\,\eta(F)\bigr]\bigr\rvert$
$\leqslant\biggl\lvert\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a,b=1}^{s}\widetilde{C}_{a,b}\bigl(\partial_{(i,a),(i,b)}\eta(F)-\partial_{(i,a),(i,b)}\eta(F-V_{ik})\bigr)\biggr]\biggr\rvert$
$+\biggl\lvert\sum_{i=1}^{d}\sum_{a,b=1}^{s}\sum_{k=1}^{m}\mathbbm{E}\bigl[W_{ik}^{2}h_{k,a}h_{k,b}\int_{0}^{1}\bigl(\partial_{(i,a),(i,b)}\eta(F-V_{ik}+sV_{ik})-\partial_{(i,a),(i,b)}\eta(F-V_{ik})\bigr)\mathrm{Leb}(ds)\bigr]\biggr\rvert$
$\leqslant\frac{1}{3}\sum_{i=1}^{d}\sum_{a,b,c=1}^{s}\mathbbm{E}\bigl\lvert\widetilde{C}_{a,b}h_{k,c}W_{ik}\bigr\rvert+\frac{1}{6}\sum_{i=1}^{d}\sum_{a,b,c=1}^{s}\mathbbm{E}\bigl\lvert W_{ik}^{3}h_{k,a}h_{k,b}h_{k,c}\bigr\rvert$
@@
where in the second to last inequality we use the bounds on the derivatives given at (2.4) (for both terms), and the final inequality uses Cauchy-Schwarz with $\mathbbm{E}[W_{ik}^{2}]=c_{2}/m$ and that $\mathbbm{E}[\lvert W_{ik}\rvert^{3}]\leqslant c_{3}/m^{3/2}$. The result now easily follows. ∎
We can use Lemma 2.1 to prove the following corollary bounding the first term of (2.2).
@@
Retaining the notation above, if $\mathbbm{E}\bigl[\lvert W_{ij}^{(L-1)}\rvert^{3}\bigr]\leqslant c_{3}^{(L-1)}n_{L-1}^{-3/2}$, then
$\mathsf{d}_{3}(F_{[d]}^{(L)}(\chi),\widetilde{F}_{[d]}^{(L)}(\chi))\leqslant\biggl(\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr)\frac{d\,s^{2}}{\sqrt{n_{L-1}}}\sum_{a=1}^{s}\mathbbm{E}\Bigl[\bigl\lvert\sigma(F_{1}^{(L-1)}(x_{a})\bigr\rvert^{3}\Bigr].$ (2.7)
@@
Proof of Corollary 2.2.
Note that from the tower property of conditional expectation, we have
$\mathsf{d}_{3}(F_{[d]}^{(L)}(\chi),\widetilde{F}_{[d]}^{(L)}(\chi))\leqslant\mathbbm{E}\Bigl[\mathsf{d}_{3}\Bigl(\mathscr{L}\bigl(F_{[d]}^{(L)}(\chi)|F^{(L-1)}(\chi)\bigr),\mathscr{L}\bigl(\widetilde{F}_{[d]}^{(L)}(\chi)|F^{(L-1)}(\chi)\bigr)\Bigr)\Bigr],$ (2.8)
and given $F^{(L-1)}(\chi)$, $F_{[d]}^{(L)}(\chi)$ and $\widetilde{F}_{[d]}^{(L)}(\chi)$ have the form of $F$ and $\widetilde{F}$ with $h_{k,a}:=\sigma(F_{k}^{(L-1)}(x_{a}))$ in Lemma 2.1. The lemma implies that
$\mathsf{d}_{3}\Bigl(\mathscr{L}\bigl(F_{[d]}^{(L)}(\chi)|F^{(L-1)}(\chi)\bigr),$ $\mathscr{L}\bigl(\widetilde{F}_{[d]}^{(L)}(\chi)|F^{(L-1)}(\chi)\bigr)\Bigr)$
$\leqslant\biggl(\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr)\frac{d}{n_{L-1}^{3/2}}\sum_{k=1}^{n_{L-1}}\biggl(\sum_{a=1}^{s}\bigl\lvert\sigma(F_{k}^{(L-1)}(x_{a})\bigr\rvert\biggr)^{3},$
$\leqslant\biggl(\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr)\frac{d\,s^{2}}{n_{L-1}^{3/2}}\sum_{k=1}^{n_{L-1}}\sum_{a=1}^{s}\bigl\lvert\sigma(F_{k}^{(L-1)}(x_{a})\bigr\rvert^{3},$
where the last inequality uses that for $x\in\mathbbm{R}^{s}$, $\|x\|_{1}\leqslant s^{2/3}\|x\|_{3}$, which follows from Hölder’s inequality.The result follows after plugging this in to (2.8) and taking expectation, noting that $F^{(L-1)}_{i}\stackrel{{\scriptstyle d}}{{=}}F^{(L-1)}_{j}$, because the weight matrices have identically-distributed rows. ∎
@@ 2.2 Error for Gaussian weights to Gaussian
Let $H\in\mathbbm{R}^{m\times s}$ be a random matrix with identically-distributed rows, independent of the matrix $\widetilde{W}\in\mathbbm{R}^{d\times m}$, which has i.i.d. centered independent Gaussian entries having variance $\mathop{\mathrm{Var}}\nolimits(\widetilde{W}_{ij})=:c_{2}/m$. Define $\widetilde{F}\in\mathbbm{R}^{d\times s}$ by
@@
$\mathbbm{E}[G_{i,a}G_{j,b}]=\delta_{ij}c_{2}\mathbbm{E}[K_{a}K_{b}],$
@@
$\mathsf{d}_{3}(\widetilde{F},G)\leqslant\frac{c_{2}\,d}{2}\sum_{a,b=1}^{s}\Bigl(\bigl\lvert\mathbbm{E}[K_{a}K_{b}]-\mathbbm{E}[H_{1,a}H_{1,b}]\bigr\rvert+\sqrt{\frac{\mathop{\mathrm{Var}}\nolimits(H_{1,a}H_{1,b})}{m}}+\sqrt{\mathrm{Cov}(H_{1,a}H_{1,b},H_{2,a}H_{2,b})}\Bigr).$
@@
[…]
@@
$\mathbbm{E}[F_{i,a}F_{j,b}|H]=\delta_{i,j}\frac{c_{2}}{m}\sum_{k=1}^{m}H_{k,a}H_{k,b}=:\delta_{i,j}C_{a,b}(H).$
@@
$\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a=1}^{s}\widetilde{F}_{i,a}\partial_{(i,a)}\eta(\widetilde{F})\Big|H\biggr]=\mathbbm{E}\biggl[\sum_{i=1}^{d}\sum_{a,b=1}^{s}C_{a,b}(H)\,\partial_{(i,a),(i,b)}\eta(\widetilde{F})\Big|H\biggr].$ (2.11)
@@
[…]
Using that the rows of $H$ are identically-distributed, we easily find
$\mathbbm{E}[C_{a,b}(H)]=c_{2}\mathbbm{E}[H_{1,a}H_{1,b}],$
@@
$\mathbbm{E}\bigl\lvert C_{a,b}(H)-\mathbbm{E}[C_{a,b}(H)]\bigr\rvert$ $\leqslant\sqrt{\mathop{\mathrm{Var}}\nolimits(C_{a,b}(H))}$
@@
Taking $H_{k,a}=\sigma(F_{k}^{(L-1)}(x_{a}))$ and $K_{a}=\sigma(G_{1}^{(L-1)}(x_{a}))$ in Lemma 2.4 (noting that $H$ has identically distributed rows as a consequence of the same property for the weight matrices) easily gives the following corollary bounding the second term of (2.2).
@@
$\mathsf{d}_{3}\bigl($ $\widetilde{F}_{[d]}^{(L)}(\chi),G_{[d]}^{(L)}(\chi)\bigr)$
$\leqslant\frac{c_{2}^{(L-1)}\,d}{2}\sum_{a,b=1}^{s}\biggl\{\Bigl\lvert\mathbbm{E}\bigl[\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\bigr]-\mathbbm{E}\bigl[\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr]\Bigr\rvert$ (2.12)
@@
Let $X,Z\in\mathbbm{R}^{d}$ be random vectors with $\mathbbm{E}[\lvert X_{i}\rvert^{q+d}]<C_{q}$ and $\mathbbm{E}[\lvert Z_{i}\rvert^{q+d}]<C_{q}$ for $i=1,\ldots,d$ and some $q>0$. Then
$\biggl\lvert\mathbbm{E}\prod_{i=1}^{d}X_{i}-\mathbbm{E}\prod_{i=1}^{d}Z_{i}\biggr\rvert\leqslant 2\sqrt{d}(4\sqrt{d}\,C_{q})^{(d-1)/(q+d-1)}\mathsf{d}_{1}(X,Z)^{q/(q+d-1)}.$
@@
Denoting the indicator function by $\mathbbm{I}$, for any coupling $(X,Z)$ of $X$ and $Z$, writing the product as a telescoping sum, and using the triangle inequality and a union-type bound gives
$\biggl\lvert\prod_{i=1}^{d}X_{i}-\prod_{i=1}^{d}Z_{i}\biggr\rvert$ $\leqslant\biggl(\sum_{j=1}^{d}\lvert X_{j}-Z_{j}\rvert\prod_{i=1}^{j-1}\lvert Z_{i}\rvert\prod_{k=j+1}^{d}\lvert X_{k}\rvert\biggr)\mathbbm{I}\Bigl[\lvert Z_{j}\rvert\leqslant T,\lvert X_{j}\rvert\leqslant T,\,\,\mathrm{for}\,\,\mathrm{all}\,\,j=1,\ldots,d\Bigr]$
$\qquad+\biggl\lvert\prod_{i=1}^{d}X_{i}-\prod_{i=1}^{d}Z_{i}\biggr\rvert\biggl(\sum_{j=1}^{d}\Bigl(\mathbbm{I}\bigl[\lvert Z_{j}\rvert\geqslant\mathrm{T}\bigr]+\mathbbm{I}\bigl[\lvert X_{j}\rvert\geqslant\mathrm{T}\bigr]\Bigr)\biggr)$
@@
$\mathbbm{E}\biggl\{\prod_{i=1}^{d}\lvert X_{i}\rvert\mathbbm{I}\bigl[\lvert Z_{j}\rvert\geqslant\mathrm{T}\bigr]\biggr\}$ $\leqslant\mathrm{T}^{-q}\mathbbm{E}\biggl\{\lvert Z_{j}\rvert^{q}\prod_{i=1}^{d}\lvert X_{i}\rvert\biggr\}$
$\leqslant\mathrm{T}^{-q}\mathbbm{E}\bigl[\lvert Z_{j}\rvert^{q+d}\bigr]^{q/(q+d)}\prod_{i=1}^{d}\mathbbm{E}\bigl[\lvert X_{i}\rvert^{q+d}\bigr]^{1/q+d}$
@@
$\biggl\lvert\mathbbm{E}\prod_{i=1}^{d}X_{i}-\mathbbm{E}\prod_{i=1}^{d}Z_{i}\biggr\rvert\leqslant\sqrt{d}\mathrm{T}^{d-1}\mathsf{d}_{1}(X,Z)+\mathrm{T}^{-q}(4d\,C_{q}).$
@@
Fix $p\in\mathbbm{N}$. If for $\ell=0,\ldots,L-2$, there are constants $c_{2p}^{(\ell)}\geqslant 1$ such that $\mathbbm{E}\bigl[(W_{ij}^{(\ell)})^{2p}\bigr]\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, then
$\max\Bigl\{\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-1)}(x))\bigr)^{2p}\bigr],\mathbbm{E}\bigl[\bigl(\sigma(G_{1}^{(L-1)}(x))\bigr)^{2p}\bigr]\Bigr\}\leqslant B_{2p}^{(L-1)}(x),$
@@
Comparing to $\sigma(0)$, using the triangle inequality for the $(2p)$-moment norm, and that $\sigma$ is Lipschitz, we find
[…]